Let f(x) = (x + 1)2 –1, x > –1
Statement – 1 : The set {x : f(x) = f–1(x)} = {0, –1}. Statement – 2 : f is a bijection.

Grade:11

3 Answers

Arun
3 years ago
Dear Jayant
 
Question is incomplete. Please check and repost the question. You can also attach an image. I will be happy to help you.
Vikas TU
3 years ago
f(x is not cont....................
Bijection - one - one function 
f(x) = x 
(x+1)^2 -1 = x 
f(x) -x = 0 
x^2 -2x + 2-1 -x = 0 
x(x+1) =0 
roots are 0,1 
statement 1 is true …...........
Vikas TU
3 years ago
Dear student 
Roots are (0, -1 )...............................................................................................................